OpenStudy (anonymous):

Consider the following pair of equations: y = –2x + 8 y = x – 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form. (5 points)

OpenStudy (anonymous):

@Godlovesme

OpenStudy (godlovesme):

r u familiar with the substitution method?

OpenStudy (anonymous):

not really that y i need help lol

OpenStudy (godlovesme):

okiee this is how u do it : make sense so far?

OpenStudy (anonymous):

yes

OpenStudy (godlovesme):

yayy lol :p so next step is solving for x can u try? :)

OpenStudy (anonymous):

ok give me a min ill try

OpenStudy (godlovesme):

alrighty :)

OpenStudy (anonymous):

x=-2+9 x+2x=9 3x=9 x=9/3 so x = 3 ???????????

OpenStudy (godlovesme):

lemme check ..

OpenStudy (godlovesme):

yes u r correct it's 3 :))) nice job *insert applause* :p

OpenStudy (godlovesme):

so to solve y, it says y=x-1 y= x-1 y= 3-1 y= 2 :3

OpenStudy (anonymous):

yes so now all i have to do is wright it out and explain how i got my answers ??

OpenStudy (godlovesme):

yes :) and dont forget to include that the ordered pair is the solution and where the lines intersect :)

OpenStudy (anonymous):

yay thank you so much

OpenStudy (godlovesme):

anytime shadow ^_^